USE OF CAVITY RING DOWN SPECTROSCOPY IN ATMOSPHERIC STUDIES

\underline{A. R. RAVISHANKARA}, S. S. BROWN, H. STARK, MATTIAS ALDENER, ANDERS PETTER-SON, and E. R. LOVEJOY, National Oceanic and Atmospheric Administration, Aeronomy Laboratory, Boulder CO,and The Cooperative Institute for Research in Environmental Sciences, and The Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO.

The Cavity ringdown spectroscopy is a very versatile tool that has found many applications in the studies of the atmosphere. I will describe the use of CRDS in laboratory studies of atmospheric chemistry and photochemistry, measurements of NO3 and N2O5 in the atmosphere, quantification of aerosol extintion, and measurements of absorption cross sections needed for climate studies. The relevant atmospheric processes will be described and the needed information will be discussed. Examples of the relevant information from our laboratory studies will be presented.